OpenStudy (anonymous):

A cylindrical container 0.8 m tall contains mercury to a certain depth, d. The rest of the cylinder is filled with water. If the pressure at the bottom of the cylinder is 1.2 atm, what is the depth d?

OpenStudy (anonymous):

Pm=pressure contribution from mercury is (13.5 x 10^3 kg/m^3)(9.8 N/kg)(d, m) Pw=pressure contribution from water is (1.0 x 10^3 kg/m^3)(9.8 N/kg)(0.8m-d) P = 1.2 atm = (1.2)(1.01 x 10^5 n/m^2) = Pm + Pw The rest is "plug and chug," solving for d.